# Learning High-Degree Parities: The Crucial Role of the Initialization

### v1

@@ Abstract
Parities have become a standard benchmark for evaluating learning algorithms. Recent works show that regular neural networks trained by gradient descent can efficiently learn degree $k$ parities on uniform inputs for constant $k$, but fail to do so when $k$ and $d-k$ grow with $d$ (here $d$ is the ambient dimension). However, the case where $k=d-O_{d}(1)$ (almost-full parities), including the degree $d$ parity (the full parity), has remained unsettled. This paper shows that for gradient descent on regular neural networks, learnability depends on the initial weight distribution. On one hand, the discrete Rademacher initialization enables efficient learning of almost-full parities, while on the other hand, its Gaussian perturbation with large enough constant standard deviation $\sigma$ prevents it. The positive result for almost-full parities is shown to hold up to $\sigma=O(d^{-1})$, pointing to questions about a sharper threshold phenomenon. Unlike statistical query (SQ) learning, where a singleton function class like the full parity is trivially learnable, our negative result applies to a fixed function and relies on an initial gradient alignment measure of potential broader relevance to neural networks learning.
@@ 2 Related Work
Previous works have studied the sample and time complexity of learning with SGD on neural networks, proposing various measures, such as: the noise sensitivity ([O’D14, ZRKB21, ABC+22, HR24]), which applies mostly to settings with i.i.d. inputs and is related to the degree of the functions, is known to be loose for strong learning ([AAM22, AAM23]); the globality degree ([ABL+24]), which generalizes the degree and sensitivity notions to non-i.i.d. settings but remains focused on weak rather than strong learning; the statistical query (SQ) dimension ([Kea98, Fel16]) and the cross-predictability ([AS20]), which are usually defined for a class of targets/distributions rather than a single distribution (in particular the full parity is efficiently SQ learnable since there is a single function); the neural tangent kernel (NTK) alignment ([JGH18, CMR12]) that are limited to the NTK framework; the information exponent ([AGJ21, BPVZ23]), generative exponent ([DPVLB24]) and leap ([AAM23]), which measure when fully connected neural networks can strongly learn target functions on i.i.d. or isotropic input distributions and sparse or single/multi-index functions. In particular, few works studied measures based on the alignment between the networks initialization and the target distribution, as in this paper. ([MNK+22, OJMDF21]) studied the label-gradient-alignment (LGA), defined as the norm of the target function in the RKHS induced by the NTK ([JGH18]) at initialization, showing its empirical relevance for predicting network performance. In contrast, we focus on a theoretical analysis, with our measure of initial gradient alignment being loss-dependent. [ACHM22] defined the initial alignment (INAL) as the maximum average correlation of any neuron with the target, providing a lower bound for functions with small INAL, though their result relies on input embedding and orbit hardness, which does not apply to almost-full parities.
@@ 3 Setting and Informal Contributions
Consider a neural network with parameters $\theta\in\mathbb{R}^{P}$ and two independent random vectors $A,H_{\sigma}\in\mathbb{R}^{P}$ with independent coordinates where $A$ is arbitrary and $H_{\sigma}$ has independent entries $H_{\sigma}\sim\mathcal{N}(0,\sigma^{2}\cdot\mathbb{I}_{P})$. We say that a neural network $\operatorname{NN}(x;\theta)$ has a $(A,\sigma)$-perturbed initialization with noise level $\sigma$ if its parameters are initialized to $\theta^{0}_{p}=A_{p}+\sqrt{\operatorname{Var}A_{p}}(H_{\sigma})_{p}$.
@@
Let $f(x)=\chi_{d}(x)$. A two-layer $\operatorname{ReLU}$ network with some $\operatorname{poly}(d)$ hidden units and Rademacher initialization, trained by GD or SGD with any batch-size with the correlation [^4] The correlation loss is defined as $L_{\rm corr}(y,\hat{y})=-y\hat{y}$. or the hinge loss, will learn $f$ to perfect accuracy in $\operatorname{poly}(d)$ steps. The result also holds for SGD with hinge loss for $\sigma$-perturbed Rademacher intitialization up to some $\sigma=O(d^{-1})$.
@@
We remark that for the squared and the correlation loss, the Gradient Alignment at initialization corresponds to the Label-Gradient-Alignment of [OJMDF21, MNK+22].
@@
Let $f(x)=\chi_{S}(x)$, for $S\subseteq[d]$ such that $|S|\geq d/2$. Assume either: 1) Gaussian initialization, or 2) $S=[d]$ (full-parity) and $\sigma$-perturbed Rademacher initialization for $\sigma>\sigma^{*}$, for some fixed $\sigma^{*}$. Then, noisy GD with correlation loss and any noise level $\tau\geq\exp(-o(d))$ in any two-layer fully connected $\operatorname{ReLU}$ network of $\exp(o(d))$ size will not achieve better accuracy than random guessing in $\exp(o(d))$ training steps.
Full versions of Theorems 1 and 3 presented in the following sections provide the following rigorous separation between Rademacher initialization and Gaussian or $\sigma$-perturbed, with $\sigma$ large enough constant: Noisy-GD for correlation loss, when applied to a two-layer fully connected $\operatorname{ReLU}$ network with some $\operatorname{poly}(d)$ hidden neurons, can learn the full parity function in $\operatorname{poly}(d)$ steps if the network is initialized with Rademacher weights. However, using Gaussian or perturbed initialization while leaving all other aspects of the algorithm unchanged requires exponential time to learn. Furthermore, the negative result is robust to details like changing hyperparameters, and the positive result is also valid for some ranges of $\sigma$-perturbed Rademacher initialization.
@@ 4 Positive Result for Rademacher Initialization
where for all $t\in\{0,\ldots,T-1\}$, $Z^{t}$ are i.i.d. $\mathcal{N}(0,\tau^{2})$, for some noise level $\tau$, and they are independent from other variables, and $B$ is the batch size. If the average over the batch size $\frac{1}{B}\sum_{s=1}^{B}\nabla_{\theta^{t}}L(\operatorname{NN}(x^{s};\allowbreak\theta^{t}),f(x^{s}))$ is replaced by the population mean $\mathbb{E}_{x\sim\mathcal{D}}\left[\nabla_{\theta^{t}}L(\operatorname{NN}(x;\theta^{t}),f(x))\right]$, we refer to the algorithm as (full batch) noisy-GD.
@@ 4.1 GD and SGD with correlation loss
The proofs are deferred to Appendix A. In the corollaries above, we have chosen convenient bias values for simplicity, but the precise values are not crucial except for ’’unlucky‘‘ choices where $\Delta$ can become too small. In particular learning should hold for random biases for most reasonable distributions. For the clipped ReLU activation, we expect (but do not prove) that the bound on the number of neurons in Corollary 2 could be improved to $n\geq\Omega(d^{2+2\lceil a/2\rceil})$ using a similar modification as in Corollary 1.
@@ 5.1 Negative Results for General Targets
5.2 Negative Results for Almost-Full Parities
@@ 7 Conclusion
In this paper, we advance the understanding of whether high degree parities can be learned using noisy-GD on standard neural networks with i.i.d. initializations. Specifically, we show that while the full parity is easily learnable with Rademacher initialization, it becomes challenging when Gaussian perturbations with large variance are introduced. This constitutes a separation between SQ algorithms and gradient descent on neural networks: the full parity is an example of a function that is trivially learnable in the statistical query (SQ) framework but difficult for noisy-GD on neural networks with most typical initializations, with the Rademacher being a special case. It raises interesting questions about a threshold where learning behavior changes based on the perturbation level $\sigma$. Additionally, we propose a novel, loss-dependent measure for assessing alignment between the initialization and the target distribution, and prove a negative result for the correlation loss that applies to general input distributions, beyond the specific case of full parity and Boolean inputs. We leave to future work strengthening of that result, e.g., to hinge loss and/or deeper architectures.
@@ 8 Acknowledgments
The authors thank Nati Srebro for a useful discussion and Ryan O‘Donnell for a helpful communication regarding binomial coefficients. Part of this work was done while EC was visiting the African Institute for Mathematical Sciences (AIMS), Rwanda. JH and DKY were supported by the Alexander von Humboldt Foundation German research chair funding and the associated DAAD project No. 57610033.
@@ A.3.2 Proof of Corollary 2
$=\frac{1}{2^{a}}\sum_{\ell=0}^{a}\binom{a}{\ell}\frac{(-1)^{d-a+c-\ell}}{2^{d-a}}\left[(d+b+2(\ell-a))\binom{d-a-2}{c-\ell-2}-(d-b-2\ell)\binom{d-a-2}{c-\ell-1}\right]\;,$ (63)
@@
The main lemma that we will need is the following bound:
@@ A.4 Almost Full Parities $d-1$ and $d-2$
The rest of Corollary 5 follows easily from Theorem 4. ∎
@@ B.1 Proof of Theorem 6
where $\xi^{t}\overset{iid}{\sim}\mathcal{N}(0,\mathbb{I}\tau^{2})$. We call $\gamma$ the learning rate and $\tau$ the noise-level of the noisy-GD algorithm used to train the network $\operatorname{NN}(x;\theta)$.
@@ B.5 Proof of Proposition 3
$\mathbb{E}_{\psi^{t}}\|\Gamma_{f}(\psi^{t})-\Gamma_{r}(\psi^{t})\|_{2}^{2}=\sum_{p=1}^{P}\mathbb{E}_{\psi^{t}}\mathbb{E}_{x}\left[\partial_{\psi^{t}_{p}}\operatorname{NN}(x;\psi^{t})\cdot f(x)\right]^{2}\;.$
@@ B.6 Proof of Lemma 6
$\partial_{W_{ij}^{(l)}}\operatorname{NN}(x;\overline{C}(\theta))=\sum_{k=1}^{N_{L-1}}C^{(L)}W_{k}^{(L)}\sigma^{\prime}(x_{k}^{(L-1)}(\theta))\cdot D_{W_{ij}^{(l)},H}\partial_{W_{ij}^{(l)}}x_{k}^{(L-1)}(\theta),\qquad l=1,...,L-1,$ (161)
@@ C.2 Lemma 7 and Lemma 8 imply Proposition 1
$\mathbb{E}_{\theta}\left(\mathbb{E}_{x}f_{a}(x)\nabla_{w_{ij}}\operatorname{NN}\right)^{2}=\mathbb{E}_{x,x^{\prime}}\left(\prod_{\ell=1}^{d-a}x_{\ell}x^{\prime}_{\ell}\right)\left(\mathbb{E}_{v_{i}}v_{i}^{2}\right)\left(\mathbb{E}_{w_{i},b_{i}}\mathds{1}(w_{i}\cdot x+b_{i}\geq 0)\mathds{1}(w_{i}\cdot x^{\prime}+b_{i}\geq 0)\right)x_{j}x^{\prime}_{j}\;.$ (178)
@@ D.4 Proof of Lemma 9
$\operatorname{GAL}_{\mathrm{hid}}$ $\leq\exp(-c\beta^{2}d)+\mathbb{E}_{x,y}\left|\mathbb{E}_{k}(-1)^{k}\mathds{1}_{\mathcal{F}}\left(\Phi(A)\Phi(w)+\phi(A)\phi(w)\sum_{\ell=0}^{\infty}H_{\ell}(A)H_{\ell}(w)\frac{\rho^{\ell+1}}{(\ell+1)!}\right)\right|$ (262)
@@
Now let us move to the left-hand side term in (273). It is sufficient to prove $1/2+\sum_{\ell<d/10}\allowbreak|a_{\ell}|(\alpha\sqrt{d})^{2\ell+1}\leq\exp(c\beta^{2}d/2)$ and this is what we are going to show. Indeed,
@@ D.5 Proof of Lemma 10
$=\mathbb{E}_{x,x^{\prime},r}\left[\prod_{i}x_{i}x^{\prime}_{i}\mathbb{E}_{g}\left[\operatorname{ReLU}(g\cdot x+\mu r\cdot x)\operatorname{ReLU}(g\cdot x^{\prime}+\mu r\cdot x^{\prime})\right]\right]:=\mathbb{E}_{x,x^{\prime},r}F(x,x^{\prime},r)\;.$ (292)
@@
$\operatorname{GAL}_{\mathrm{hid}}$ $\leq\exp(-c\beta^{2}d)+\mathbb{E}_{x,y}\Bigg{|}\mathbb{E}_{k}(-1)^{k}\mathds{1}_{\mathcal{F}}\Bigg{(}R(A)R(w)+\Phi(A)\Phi(w)\rho$ (299)
$\qquad\qquad\qquad\qquad\qquad\qquad+\phi(A)\phi(w)\sum_{\ell=0}^{\infty}\frac{H_{\ell}(A)H_{l}(w)}{(\ell+2)!}\rho^{\ell+2}\Bigg{)}\Bigg{|}$ (300)

### v3

@@ Abstract
Parities have become a standard benchmark for evaluating learning algorithms. Recent works show that regular neural networks trained by gradient descent can efficiently learn degree $k$ parities on uniform inputs for constant $k$, but fail to do so when $k$ and $d-k$ grow with $d$ (here $d$ is the ambient dimension). However, the case where $k=d-O_{d}(1)$, including the degree $d$ parity (the full parity), has remained unsettled. This paper shows that for gradient descent on regular neural networks, learnability depends on the initial weight distribution. On one hand, the discrete Rademacher initialization enables efficient learning of full parities, while on the other hand, its Gaussian perturbation with large enough constant standard deviation $\sigma$ prevents it. The positive result for full parity is shown to hold up to $\sigma=O(d^{-1})$, pointing to questions about a sharper threshold phenomenon. Unlike statistical query (SQ) learning, where a singleton function class like the full parity is trivially learnable, our negative result applies to a fixed function and relies on an initial gradient alignment measure of potential broader relevance to neural networks learning.
@@ 1 Introduction
Parity functions are a well-known class of challenging problems for differentiable learning models, where the task is to determine the parity of bits belonging to an unknown subset of input coordinates. Due to their inherent non-linearity and extreme sensitivity to small input changes, parity functions often serve as a challenging benchmark for evaluating and comparing learning algorithms, including gradient descent on neural networks ([AS20, DM20]). For instance, they have been used for showing the advantages of using convolutional architectures over fully connected ones ([MSS20]), the superiority of differentiable models compared to kernel methods ([AKM+21]), and the efficacy of curriculum learning in contrast to standard training ([ACL24, CM23]).
@@ 2 Related Work
Previous works have studied the sample and time complexity of learning with SGD on neural networks, proposing various measures, such as: the noise sensitivity ([O’D14, ZRKB21, ABC+22, HR24]), which applies mostly to settings with i.i.d. inputs and is related to the degree of the functions, is known to be loose for strong learning ([AAM22, AAM23]); the globality degree ([ABL+24]), which generalizes the degree and sensitivity notions to non-i.i.d. settings but remains focused on weak rather than strong learning; the statistical query (SQ) dimension ([Kea98, Fel16]) and the cross-predictability ([AS20]), which are usually defined for a class of targets/distributions rather than a single distribution (in particular the full parity is efficiently SQ learnable since there is a single function); the neural tangent kernel (NTK) alignment ([JGH18, CMR12]) that are limited to the NTK framework; the information exponent ([AGJ21, BPVZ23]), generative exponent ([DPVLB24, DTA+24]), leap ([AAM23]) and Approximate Message Passing (AMP) complexity ([TDD+24]), which measure when fully connected neural networks can strongly learn target functions on i.i.d. or isotropic input distributions and sparse or single/multi-index functions. In particular, few works studied measures based on the alignment between the networks initialization and the target distribution, as in this paper. ([MNK+22, OJMDF21]) studied the label-gradient-alignment (LGA), defined as the norm of the target function in the RKHS induced by the NTK ([JGH18]) at initialization, showing its empirical relevance for predicting network performance. In contrast, we focus on a theoretical analysis, with our measure of initial gradient alignment being loss-dependent. [ACHM22] defined the initial alignment (INAL) as the maximum average correlation of any neuron with the target, providing a lower bound for functions with small INAL, though their result relies on input embedding and orbit hardness, which does not apply to almost-full parities.
@@ 3 Setting and Informal Contributions
Consider a neural network with parameters $\theta\in\mathbb{R}^{P}$ and two independent random vectors $A,H_{\sigma}\in\mathbb{R}^{P}$ with independent coordinates where $A$ is arbitrary and $H_{\sigma}$ has independent entries $(H_{\sigma})_{p}\sim\mathcal{N}(0,\sigma^{2}\cdot\mathbb{I}_{P})$. We say that a neural network $\operatorname{NN}(x;\theta)$ has a $(A,\sigma)$-perturbed initialization with noise level $\sigma$ if its parameters are initialized to $\theta^{0}_{p}=A_{p}+\sqrt{\operatorname{Var}A_{p}}(H_{\sigma})_{p}$.
@@
Let $f(x)=\chi_{d}(x)$. A two-layer $\operatorname{ReLU}$ network with some $\operatorname{poly}(d)$ hidden units and $\sigma$-perturbed Rademacher initialization with $\sigma=O(d^{-1})$, trained by GD or SGD with any batch-size with the correlation [^4] The correlation loss is defined as $L_{\rm corr}(y,\hat{y})=-y\hat{y}$. or the hinge loss, will learn $f$ to perfect accuracy in $\operatorname{poly}(d)$ steps.
@@
We remark that for the squared and the correlation loss, the Gradient Alignment at initialization corresponds to the Label-Gradient-Alignment (LGA) of [OJMDF21, MNK+22], thus our GAL generalizes LGA to other losses.
@@
Let $f(x)=\chi_{S}(x)$, for $S\subseteq[d]$ such that $|S|\geq d/2$. Noisy-GD with correlation loss and any noise level $\tau=\Omega(1/\operatorname{poly}(d))$ on any two-layer fully connected $\operatorname{ReLU}$ network of $\operatorname{poly}(d)$ size, initialized with Gaussian initialization will not achieve accuracy better than random guessing in $\operatorname{poly}(d)$ training steps.
We expect Theorem 3 to hold also in case of $\sigma$-perturbed Rademacher initialization for $\sigma>\sigma^{*}$ for some fixed $\sigma^{*}>0$. To that end in Section 5.2.2 we prove the gradient alignment bound for the hidden layer weights in the perturbed case. Together with a similar bound for the output layer weights (which we omit from this version of the paper) that would give the statement of Theorem 3 also for the $\sigma$-perturbed initialization, with $\sigma>\sigma^{*}$.
Full versions of Theorems 1 and 3 presented in the following sections provide the following rigorous separation between Rademacher and Gaussian initializations: Noisy-GD for correlation loss, when applied to a two-layer fully connected $\operatorname{ReLU}$ network with some $\operatorname{poly}(d)$ hidden neurons, can learn the full parity function in $\operatorname{poly}(d)$ steps if the network is initialized with Rademacher weights. However, using Gaussian initialization while leaving all other aspects of the algorithm unchanged requires exponential time to learn. Furthermore, the negative result is robust to details like changing hyperparameters, and as discussed above, both positive and negative results are also valid for some ranges of $\sigma$-perturbed Rademacher initializations.
@@ 4 Positive Result for Rademacher Initialization
where for all $t\in\{0,\ldots,T-1\}$, $Z^{t}$ are i.i.d. $\mathcal{N}(0,\tau^{2})$, for some noise level $\tau$, and they are independent from other variables, and $B$ is the batch size. If the average over the batch size $\frac{1}{B}\sum_{s=1}^{B}\nabla_{\theta^{t}}L(\operatorname{NN}(x^{s};\theta^{t}),f(x^{s}))$ is replaced by the population mean $\mathbb{E}_{x\sim\mathcal{D}}\left[\nabla_{\theta^{t}}L(\operatorname{NN}(x;\theta^{t}),f(x))\right]$, we refer to the algorithm as (full batch) noisy-GD.
@@ 4.1 GD and SGD with correlation loss
In the corollaries above, we have chosen convenient bias values for simplicity, but the precise values are not crucial except for ’’unlucky‘‘ choices where $\Delta$ can become too small. In particular learning should hold for random biases for most reasonable distributions. For the clipped ReLU activation, we expect (but do not prove) that the bound on the number of neurons in Corollary 2 could be improved to $n\geq\Omega(d^{2+2\lceil a/2\rceil})$ using a similar modification as in Corollary 1.
@@ 5.1 Negative Results for General Targets
5.2 Negative Results for High-Degree Parities
@@ 7 Conclusion
In this paper, we advance the understanding of whether high degree parities can be learned using noisy-GD on standard neural networks with i.i.d. initializations. Specifically, we show that while the full parity is easily learnable with Rademacher initialization, it becomes challenging when Gaussian perturbations with large variance are introduced. This constitutes a separation between SQ algorithms and gradient descent on neural networks: the full parity is an example of a function that is trivially learnable in the statistical query (SQ) framework but difficult for noisy-GD on neural networks with most typical initializations, with the Rademacher being a special case. It raises interesting questions about a threshold where learning behavior changes based on the perturbation level $\sigma$; result, e.g., to hinge loss and/or deeper architectures.
Additionally, we propose a novel, loss-dependent measure for assessing alignment between the initialization and the target distribution, and prove a negative result for the correlation loss that applies to general input distributions, beyond the specific case of full parity and Boolean inputs. We leave to future work investigating if this technique can be applied more generally, especially in other scenarios where the dynamics remain stuck near initialization for a significant time (hardness of weak-learning). For example, such behavior is plausible for targets presenting symmetries and requiring some level of ’logical reasoning’ (e.g. arithmetic, graphs, syllogisms), for which parities are a simple model. Similarly, we leave open strengthening of the negative result, e.g., to hinge loss and/or deeper architectures.
@@ 8 Acknowledgments
The authors thank Nati Srebro for a useful discussion and Ryan O‘Donnell for a helpful communication regarding binomial coefficients. Part of this work was done while EC was visiting the African Institute for Mathematical Sciences (AIMS), Rwanda. JH and DKY were supported by the Alexander von Humboldt Foundation German research chair funding and the associated DAAD project No. 57610033. EC was partly supported by the NSF DMS-2031883 and the Bush Faculty Fellowship ONR-N00014-20-1-2826.
@@ A.3.2 Proof of Corollary 2
$=\frac{1}{2^{a}}\sum_{\ell=0}^{a}\binom{a}{\ell}\frac{(-1)^{d-a+c-\ell}}{2^{d-a}}\Bigg{[}(d+b+2(\ell-a))\binom{d-a-2}{c-\ell-2}$ (63)
$\qquad\qquad-(d-b-2\ell)\binom{d-a-2}{c-\ell-1}\Bigg{]}\;,$ (64)
@@
The main lemma that we will need is the following combinatorial bound:
@@ A.4 Almost Full Parities $d-1$ and $d-2$
The rest of Corollary 5 follows easily from Theorem 4.
@@ B.1 Proof of Theorem 6
where $\xi^{t}\overset{iid}{\sim}\mathcal{N}(0,\mathbb{I}\tau^{2})$ We call $\gamma$ the learning rate and $\tau$ the noise-level of the noisy-GD algorithm used to train the network $\operatorname{NN}(x;\theta)$.
@@ B.5 Proof of Proposition 3
$\mathbb{E}_{\psi^{t}}\|\Gamma_{f}(\psi^{t})-\Gamma_{r}(\psi^{t})\|_{2}^{2}=\sum_{p=1}^{P}\mathbb{E}_{\psi^{t}}\mathbb{E}_{x}\left[\partial_{\psi^{t}_{p}}\operatorname{NN}(x;\psi^{t})\cdot f(x)\right]^{2}$
@@ B.6 Proof of Lemma 6
$\partial_{W_{ij}^{(l)}}\operatorname{NN}(x;\overline{C}(\theta))=\sum_{k=1}^{N_{L-1}}C^{(L)}W_{k}^{(L)}\sigma^{\prime}(x_{k}^{(L-1)}(\theta))\cdot D_{W_{ij}^{(l)},H}\partial_{W_{ij}^{(l)}}x_{k}^{(L-1)}(\theta),$ (163)
$\qquad\qquad\qquad\qquad\qquad l=1,...,L-1,$ (164)
@@ C.2 Lemma 7 and Lemma 8 imply Proposition 1
$\mathbb{E}_{\theta}\left(\mathbb{E}_{x}f_{a}(x)\nabla_{w_{ij}}\operatorname{NN}\right)^{2}$ $=\mathbb{E}_{x,x^{\prime}}\left(\prod_{\ell=1}^{d-a}x_{\ell}x^{\prime}_{\ell}\right)\left(\mathbb{E}_{v_{i}}v_{i}^{2}\right)$ (181)
$\qquad\qquad\cdot\left(\mathbb{E}_{w_{i},b_{i}}\mathds{1}(w_{i}\cdot x+b_{i}\geq 0)\mathds{1}(w_{i}\cdot x^{\prime}+b_{i}\geq 0)\right)x_{j}x^{\prime}_{j}\;.$ (182)
@@ D.4 Proof of Lemma 9
$\operatorname{GAL}_{\mathrm{hid}}$ $\leq\exp(-c\beta^{2}d)$ (266)
$\qquad+\mathbb{E}_{x,y}\left|\mathbb{E}_{k}(-1)^{k}\mathds{1}_{\mathcal{F}}\left(\Phi(A)\Phi(w)+\phi(A)\phi(w)\sum_{\ell=0}^{\infty}H_{\ell}(A)H_{\ell}(w)\frac{\rho^{\ell+1}}{(\ell+1)!}\right)\right|$ (267)
@@
Now let us move to the left-hand side term in (278). It is sufficient to prove $1/2+\sum_{\ell<d/10}|a_{\ell}|(\alpha\sqrt{d})^{2\ell+1}\leq\exp(c\beta^{2}d/2)$ and this is what we are going to show. Indeed,
@@ D.5 Proof of Lemma 10
$=\mathbb{E}_{x,x^{\prime},r}\left[\prod_{i}x_{i}x^{\prime}_{i}\mathbb{E}_{g}\left[\operatorname{ReLU}(g\cdot x+\mu r\cdot x)\operatorname{ReLU}(g\cdot x^{\prime}+\mu r\cdot x^{\prime})\right]\right]$ (297)
$:=\mathbb{E}_{x,x^{\prime},r}F(x,x^{\prime},r)\;.$ (298)
@@
$\operatorname{GAL}_{\mathrm{hid}}$ $\leq\exp(-c\beta^{2}d)+\mathbb{E}_{x,y}\Bigg{|}\mathbb{E}_{k}(-1)^{k}\mathds{1}_{\mathcal{F}}\cdot$ (305)
$\qquad\left(R(A)R(w)+\Phi(A)\Phi(w)\rho+\phi(A)\phi(w)\sum_{\ell=0}^{\infty}\frac{H_{\ell}(A)H_{l}(w)}{(\ell+2)!}\rho^{\ell+2}\right)\Bigg{|}$ (306)